PURE ROTATIONAL SPECTROSCOPY OF TRANSITION METAL NITRIDES: MILLIMETER-WAVE STUDIES OF FeN (X^{2}Δ_{i}) AND CrN (X^{4}Σ^{−})

\underline{P. M. SHERIDAN}, M. A. BREWSTER, L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of FeN in its X^{2}Δ_{i} electronic ground state has been observed in the region of 190 - 530 GHz using direct absorption techniques. The spectrum of CrN has been recorded in its X^{4}Σ^{−} state as well. These nitrides were formed by reacting metal vapor, produced by a Broida-type oven, with either N_{2} or NH_{3} under d.c. discharge conditions. For FeN, eight rotational transitions were measured in the Ω = 5/2 spin orbit component, where no evidence of Λ-type doubling was observed. For CrN, five rotational transitions were recorded in which all fine structure components were resolved. These data have been analyzed and rotational and certain fine-structure parameters have been determined. Bonding in metal nitride species is also discussed.